if psinthita=q then provethat underroot (p2+p2)​ ×tan theta=q

Grade:8

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Consider p = sqrt 2

Theta = 45

Then q = 1

Lets verify the condition
2 (p*p ) * tan theta = 4*1= 4

q = 1

Lhs = rhs

Kindly check the question and re-post with the correct question

Regards